The difference between the squares of two numbers is 5. Three times the square of the first number increased by the square of the second number is 31. Find the numbers.
Answer by Boreal(15235)

numbers are are x and y
x^2-y^2=5
3x^2+y^2=31
Add them
4x^2=36, the y^2 disappear
x^2=9
x=3 and -3
check both
x=3, y=2 or -2
x=-3, y=2 or -2
(3, 2) (3, -2) (-3, 2) (-3, -2)